OpenStudy (anonymous):

a boat can go 51mph in still water. It takes as long to go 640 miles upstream as it does to go downstream 720 miles. How fast is the current going?

OpenStudy (anonymous):

@welshfella do you know?

OpenStudy (welshfella):

Going downstream the boats speed is 51 + v where v is the speed of the current Going upstream its speed will by 51 - v

OpenStudy (welshfella):

Going downstream speed = distances / time 51 + v = 720 / t upstream 51 - v = 640 / t

OpenStudy (anonymous):

@welshfella they don't give us the time though.

OpenStudy (welshfella):

so downstream t = 720 / (51+ v) uostream t = 640 / (51 - v) as the times are the same 720 / (51 + v) = 640 / (51 - v) solve for v

OpenStudy (welshfella):